OpenStudy (legomyego180):

Integrate h(x, y) = (x + y) i + (y2 - x) j over the closed curve that begins at (-7, 0), goes along the x-axis to (7, 0) and returns to (-7, 0) by the upper part of the circle x2 + y2 = 49

OpenStudy (legomyego180):

Integrate h(x, y) = (x + y) i + (y^2 - x) j over the closed curve that begins at (-7, 0), goes along the x-axis to (7, 0) and returns to (-7, 0) by the upper part of the circle x^2 + y^2 = 49

OpenStudy (legomyego180):

(This is using greens theorum)

zepdrix (zepdrix):

P = x+y Py = 1 Q = y^2-x Qx = -1 \[\large\rm \oint_c (x+y)dx+(y^2-x)dy\quad=\quad \int\limits\int\limits_D-1-1~dA\]We get something like that, yes? Are you having trouble with the boundaries?

zepdrix (zepdrix):



zepdrix (zepdrix):